A rectangular air mattress of length 2m width 0.5m and thickness 0.08m. what is the maximum mass of a man who lying on mattress can float on water if the mass of the mattress is 2kg. What is the density of the mattress?
 

Grade:11

2 Answers

harshin nechooli
5 years ago
78 kg   here the total mass of the matrix + that of men will be equal to the maximum buoyancy force that the matrix can provide and then you can you create both who's this that his mass of Man Plus mattress into G equal to thousand x volume of matrix and that will give you the total marks of matrix plus man t to be 80 then the mass of man is 78
Loli Patel
3 years ago
i}  (Mass of man + mass of matress) g = density of water × volume of matress ×g
 
Let m= mass of man
=> m + 2 = 103 × 2×0.08×0.5
=> m=78kg
 
ii} let d = density of matress 
 
=> 78 + d× 2×0.08×0.5 = 103 ×2×0.08×0.5
=> 78 + 0.08d = 80
=> 0.08d = 80-78
=> d = 2/0.08
=> d= 25 kg/m³
 
Thank u..😊